OpenStudy (anonymous):

What is the value of n in the equation 1/2(n – 4) – 3 = 3 – (2n + 3)?

OpenStudy (anonymous):

@TheSmartOne Last one

TheSmartOne (thesmartone):

\[\frac{ n-4 }{ 2 }-3=3-2n-3\]

umerlodhi (umerlodhi):

the smart one i tagged you in my question please help

TheSmartOne (thesmartone):

\[\frac{ n-4 }{ 2}-3=-2n\]

TheSmartOne (thesmartone):

\[\frac{ n-4 }{ 2 }=-2n+3\]

TheSmartOne (thesmartone):

\[n-4=-4n+6\]

TheSmartOne (thesmartone):

\[5n=10\]

TheSmartOne (thesmartone):

n=2

TheSmartOne (thesmartone):

Did you understand that @justsmile531

OpenStudy (anonymous):

Yes I did. Thank you

TheSmartOne (thesmartone):

no problem...